a)     A simple closed economy, with an mpc equal to 0.75. The government has passed a balanced budget amendment. The economy goes into a recession, so the government increases government spending by 40 million to try to expand the economy

 

                   i.           What is the net effect on output from these two policies? Was there any expansionary effect? 


1
Expert's answer
2021-07-05T08:59:32-0400

The change in output from tax hike can be calculated with the help of tax multiplier.

Tax multiplier is "=\frac{-MPC}{1−MPC}"


Tax multiplier is"=\frac{-0.75}{1−0.75}"

Tax multiplier is"=\frac{-0.75}{0.25}"

Tax multiplier is"-3"

"∆Y= −3×∆T\\∆Y=−3×40 million\\∆Y= −120 million"

So, change in output from tax hike is -120 million.


The net effect of the two policies can be calculated by the sum of change in output due to two policies.

"∆Y(i+ii) = 160 million −120 million\\∆Y(i+ii) = 40 million"

Since the net effect is positive thus there was expansionary effect as output expanded by 40 million.